Let f and g be the functions defined by f(x)= 2x+3 g(x)= 3x+2 then composition of f and g is

1
Expert's answer
2022-06-14T13:12:20-0400

The composition of two functions f and g is the new function we get by performing g first, and

then performing f:

"f\circ g(x)=f(g(x))=f(3x+2)=\\
=2\cdot(3x+2)+3=2\cdot3x+2\cdot2+3=\\
=6x+7."